OpenStudy (anonymous):

√9x^2 + x√100x - √36x^3

OpenStudy (espex):

What if you rewrote it like this: \[\sqrt{3^{2}x^{2}}+x\sqrt{10^{2}x}-\sqrt{6^{2}x^{2+1}}\]